OpenStudy (anonymous):

A man stands on the roof of a building that is 50.0m tall and throws a rock with a velocity of 35 m/s at an an? a man stands on the roof of a building that is 50.0m tall and throws a rock with a velocity of 35 m/s at an angle of 50 above the horizontal: calculate a) the maximum height above the building b) the two times at which the rock is 20 meters above the building c) the velocity of the rock just before it strikes the ground d) the horizontal distance from the base of the building to the point where the rock strikes the ground.

OpenStudy (anonymous):

I'm sure you have studied projectile motion basics.

OpenStudy (anonymous):

Chose proper axes and write equations along them. patiently proceed , you'll be able to solve it. Ask any doubts you have.

OpenStudy (anonymous):

yes but all i know is 35sin50. average velocity is 1/2 vi +vf = 1/2 (35sin50) time is 0 x 35 x sin 50 + -9.8 (gravity) = 35sin50 +9.8

OpenStudy (anonymous):

have you read eqns. like v = u + at , y = ut +.5at^2 ?

OpenStudy (anonymous):

the answer is a) 35.9 m b) 17.856 m/s c) 41.44 m/s d) 153.545 m

OpenStudy (anonymous):

sorry b) 0.895 sekon

OpenStudy (anonymous):

how did you get that? step by step?

OpenStudy (anonymous):